OpenStudy (anonymous):

the solution to the equation 3.4t=17.68 is t = 5.8 true of false

OpenStudy (cggurumanjunath):

false

OpenStudy (cggurumanjunath):

(3.4t)/3.4=17.68/3.4 t=5.2 not 5.8

OpenStudy (cggurumanjunath):

is it clear @mirandaT ?